A glacier moves with a speed of 45 nm/s.
How many years would it take for the
glacier to move 0.87 km? Assume there are
365 days in a year.
Answer in units of y

Answer by colliefan(242)

I would solve this by using the "unit factor" method. It is really really helpful in chemistry and physics and all sorts of real world problems.
Basically, if you multiply something by 1, you continue to have the same value. If you have a ratio of two things that are the same like 1 foot/12 inches, that ratio is equal to 1 and so you change the units that the value is represented in, but not the value itself. If you keep multiplying by 1 in the form of ratios that equal 1, you change step by step to the value that you want.
We want to change 0.87 km to years. Start with what you know.
0.87 km * 1000 meters/1 km * 10^9 nm/1 m = 8.7*10^12 nm
If you write this out as ratios, cancel out units in the numerator and denominator to get to where you want. OK, now we've gotten as far as the nm that we need because we know that it moves 45 nm/s. We just keep going with that unit factor.
8.7*10^12 nm * 1 second/45 nm = 1.9 *10^11 seconds
The nm cancel and you're left with seconds. Now you need to convert to years. Do this with more unit factors.
1.9 *10^11 seconds * 1 min/60 sec * 1 hour/60 min * 1 day/24 hours * 1 year/365 days

This converts the units to years.
Calculate: 1.9*10^11*1/60*1/60*1/24*1/365 = to come up with the number
I get 6024.86 years.
You can search online for information about the unit factor method. Ask your teacher or a nice chemistry teacher to show it to you. It is often taught in chemistry because it is so helpful there.